((This is the last thing I'll ask please just answer only these! ))

A) How many significant digits do the following measured quantities have?

i) 2.83 cm

iii) 14.0 g

v) 0.02 mL

vii) 2.350 x 10–2 L

ix) 3 fingers

ii) 36.77 mm

iv) 0.0033 kg

vi) 0.2410 km

viii) 1.00009 L

x) 0.0056040 g

B) How many significant digits do the following measured quantities have?

i) 83.25 – 0.1075

ii) 4.02 + 0.001

iii) 0.2983 + 1.52

C) How many significant digits do the following measured quantities have?

i) 7.255 ÷ 81.334

ii) 1.142 x 0.002

iii) 31.22 x 9.8

D) Solve the following:

i) 6.12 x 3.734 + 16.1 ÷ 2.3

ii) 0.0030 + 0.02

iii) 1.70 x 103 + 1.34 x 105

iv) (33.4 + 112.7 + 0.032)/(6.487)

E) Convert these measurements:

i) 1.0 cm = ______ m

ii) 0.0390 kg = ______ g

iii) 1.7 m = _____ mm


((Sorry for the trouble but I would just like this answered and not the previous comments))

-Cassi <3


1
Expert's answer
2021-11-15T10:50:37-0500

A) i) 3 significant digits

iii) 2 significant digits

v) 1 significant digit

vii) 3 significant digits

ix) 3 significant digits

ii) 4 significant digits

iv) 2 significant digits

vi) 3 significant digits

viii) 6 significant digits

x) 6 significant digits

B) i) 83.25 – 0.1075 = 83,1425 (6 significant digits)

ii) 4.02 + 0.001 = 4.021 (4 significant digits)

iii) 0.2983 + 1.52 = 1,8183 (5 significant digits)

C) i) 7.255 ÷ 81.334 = 0.0892 (4 significant digits)

ii) 1.142 x 0.002 = 0,002284 (6 significant digits)

iii) 31.22 x 9.8 = 305,96 (5 significant digits)

D) i) 6.12 x 3.734 + 16.1 ÷ 2.3 = 29,852

ii) 0.0030 + 0.02 = 0.023

iii) 1.70 x 103 + 1.34 x 105 = 315,8

iv) (33.4 + 112.7 + 0.032)/(6.487) = 22,527

E) i) 1.0 cm = 0.01 m

ii) 0.0390 kg = 39 g

iii) 1.7 m = 1700 mm